the length of a rectangle is 4 inch less than twice the width find the Length and width if the area is 720
Answer by mananth(16946)

let width be x
length = 2x-4
..
Area = L*W
x(2x-4)=720
2x^2-4x=720
/2
x^2-2x=360
x^2-2x-360=0
x^2-20x+18x-260=0
x(x-20)+18(x-20)=0
(x-20)(x+18)=0
x= 20 in. the with.
length = 2x-4
=40-4
=36 in.
m.ananth@hotmail.ca